# Enumerating log rational curves on some toric varieties

### v2

@@ 2.4. Incidence loci
on ${\mathbb{P}}(\mathcal{E}_{1})$. The equations (2) are the degeneracy loci of maps of line bundles
@@
where $\mathcal{O}_{Q}(-1)$ denotes the universal sub-line bundle on $Q_{\Gamma}(X_{r,s,a})$.
@@ 3. Transversality
is of expected dimension 0. The degree of $V$ (as a 0-cycle) is therefore a candidate for the value of $\mathrm{logTev}^{X_{r,s,a}}_{\Gamma}$. In order to guarantee that $\deg(V)=\mathrm{logTev}^{X_{r,s,a}}_{\Gamma}$ (that is, the intersection number $\deg(V)$ is enumerative), we want $V$ to be reduced of dimension 0, and in addition supported in the locus $\mathcal{M}_{\Gamma}(X)\subset Q_{\Gamma}(X)$.
@@ 3.1. Transversality for maps
where the last map is projection. If $n=\frac{m}{r+s}+1$, then this composition is a map between smooth varieties of the same dimension, so the first part of (i) follows from generic smoothness. To see that
@@ 3.2. Twisting
Denote by $f^{\prime}$ the value of $f$ after these steps are complete, and by $\Gamma^{\prime}$ the new tangency data. Output $f^{\prime}\in Q^{\neq 0}_{\Gamma^{\prime}}(X_{r,s,a})$.
@@
The first inequality is an equality if and only if no twisting is applied at points $q\neq p_{i}$, and the second inequality is an equality if and only if no twisting is applied at the $p_{i}$. However, because the given $p_{i},x_{i}$ are chosen generally, $\mathcal{V}$ must dominate $({\mathbb{P}}^{1})^{n}\times X^{n}$, so $\dim(\mathcal{V}^{\prime})\geq\dim(({\mathbb{P}}^{1})^{n}\times X^{n})$. Therefore, equality must hold everywhere. This means that no twisting was applied at all, which is to say that $f\in\mathcal{M}^{\text{bpf}}_{\Gamma}(X_{r,s,a})$. ∎
@@ 3.3. Vanishing sections
Combining with the fact that $m_{0}\geq\ell$ and $m=(n-1)(r+s)$ gives
@@ 4.1. Vanishing
which is equivalent to $m_{\ell}\leq n-1$. If instead $r=1$, then for $\ell=1,2$, the $(s+1)$-tuple of sections $[g_{0}g_{\ell}^{a}:g_{3}:\cdots:g_{s+2}]$ defines a (log) map to ${\mathbb{P}}^{1}\to{\mathbb{P}}^{s}$ satisfying $n$ incidence conditions at general points. We find similarly, repeating the proof of Proposition 3.1.1 for $X={\mathbb{P}}^{s}$, that $m_{\ell}\leq n-1$.
Next, consider $\ell\in\{r+2,\ldots,r+s+1\}$. If $s>1$, then the sections $g_{0},\ldots,\widehat{g_{\ell}}\ldots,g_{r+s+1}$ define a point of $f^{\prime}\in\mathcal{M}_{\Gamma^{\prime}}(X_{r,s-1,a})$ satisfying $n$ general incidence conditions. Arguing as above gives $m_{\ell}\leq n-1$ once more. If $s=1$, then we consider instead the map $g:{\mathbb{P}}^{1}\to{\mathbb{P}}^{r}$ defined by $[g_{1}:\cdots:g_{r+1}]$. We obtain in this case the stronger inequality $m_{0}+m_{r+2}\leq n-1$.
Finally, we prove the inequality $m_{r+2}+\cdots+m_{r+s+1}\geq(s-1)(n-1)$. There is nothing to check if $s=1$, so we assume that $s>1$, and consider the map $h:{\mathbb{P}}^{1}\to{\mathbb{P}}^{s-1}$ defined by $[g_{r+2}:\cdots:g_{r+s+1}]$, which satisfies $n$ general incidence conditions. Therefore, we must have
@@ 4.2. Integral computation
parametrizes choices of the divisors $D_{j}$, and $\nu:{\mathbb{P}}^{1}\times\mathcal{B}\longrightarrow Y$ is the projection. The universal divisors are denoted $\mathcal{D}_{j}\subset{\mathbb{P}}^{1}\times\mathcal{B}$. We have the rank $r+1$ vector bundle
@@
For any $p\in{\mathbb{P}}^{1}$ and $x\in X^{\circ}$, the degeneracy locus description of $V(p,x)$ of §2.4 gives
@@
$=\zeta_{1}^{r}\zeta_{2}^{s}.$
@@
$\int_{{\mathbb{P}}(\mathcal{E}_{2})}\prod_{i=1}^{n}[V(p_{i},x_{i})]^{n}$ $=\int_{\mathcal{Y}}\sum_{k_{0}+k_{r+2}+\cdots k_{r+s+1}=s(n-1)}\sum_{\ell=0}^{k_{0}}\binom{k_{0}}{\ell}a^{k_{0}-\ell}\mathcal{S}_{r(n-1)+k_{0}-\ell}(\mathcal{E}_{1})\mathrm{H}_{0}^{\ell}\mathrm{H}_{r+2}^{k_{r+2}}\cdots\mathrm{H}_{r+s+1}^{k_{r+s+1}}$
$=\int_{\mathcal{Y}}\sum_{k_{0}+k_{r+2}+\cdots k_{r+s+1}=s(n-1)}\sum_{\ell=0}^{k_{0}}\binom{k_{0}}{\ell}a^{k_{0}-\ell}\sum_{k_{1}+\cdots k_{r+1}=r(n-1)+k_{0}-\ell}\mathrm{H}_{0}^{\ell}\prod_{j=1}^{r+s+1}\mathrm{H}_{j}^{k_{j}}.$
@@
because $n=\frac{m}{r+s}+1$. Note that this is only valid if $\ell=m_{0}\leq k_{0}$, because $\ell$ ranges between 00 and $k_{0}$ in the integrand. Thus, if $m_{0}>k_{0}$, then the integral is zero.
@@ 4.3. Specializations
Theorem 4.2.1 recovers [CIL24, Theorem 7] in the case of Hirzebruch surfaces, $r=s=1$. Our $m_{0},m_{1},m_{2},m_{3}$ are the $|\mu_{2}|,|\mu_{1}|,|\mu_{3}|,|\mu_{4}|$, respectively, appearing in [CIL24, Theorem 7]. If $s=1$ and $r\geq 1$ is arbitrary, then Theorem 4.2.1 proves [CIL24, Conjecture 14], and also makes precise exactly when the logarithmic Tevelev degree is zero. Taking $s$ arbitrary gives a more general statement.
The variety $X_{r,s,a}$ is Fano if and only if $a\leq s$. When in addition $a\geq 2$, the fixed-domain Gromov-Witten invariants of $X_{r,s,a}$ are shown not to be enumerative in [BLL+24, §3.2], even in genus 0 for curve classes of large anti-canonical degree, disproving the main conjecture of [LP25]. In contrast, the corresponding logarithmic invariants $\mathrm{logTev}^{X_{r,s,a}}_{\Gamma}$, with $\mu_{j}=(1)^{m_{j}}$, are enumerative in genus 0, and are determined by Theorem 4.2.1. The logarithmic setting is broadly understood to be more likely to yield enumerative counts of curves.
When $a=1$, we have $X_{r,s,a}\cong\mathrm{Bl}_{{\mathbb{P}}^{s-1}}{\mathbb{P}}^{r+s}$. Taking all of the $m_{j,v}$ equal to 1, substituting $c=d$ and $b=d-k$, and dividing by the factors of $m_{j}!$ that count orderings of the intersection points with the boundary gives that
@@
When $s=1$, this recovers the $\ell=1$ case of [CL23a, Theorem 12] for the class $\beta$ of degree $d$ curves (against the hyperplane class) meeting the exceptional divisor with degree $k$. Without the logarithmic conditions, the $s>1$ case is new, but should also be no more difficult (even in higher genus, with some additional positivity assumptions on $\beta$) using the same methods of [CL23a].
@@
Without tangency conditions, this specializes simply to $\mathrm{Tev}^{{\mathbb{P}}^{r}}_{0,n,d}=1$ [FL23, Theorem 1.1].
@@ 5. Blow-ups of ${\mathbb{P}}^{2}$
The value of $\mathrm{logTev}^{X}_{\Gamma}$ when $X$ is the blow-up of ${\mathbb{P}}^{r}$ at $r$ points has also been predicted in [CIL24, Conjecture 15]. In this section, we show that the prediction does not always hold. For notational simplicity, we take $r=2$, though the arguments hold more generally.
@@
This is similar to the situation of $\mathrm{logTev}^{X_{r,s,a}}_{\Gamma}$, where we were able to show that, in the remaining cases, the logarithmic Tevelev degree vanishes. In the setting of blow-ups of ${\mathbb{P}}^{r}$, the range in which $\mathrm{logTev}^{X}_{\Gamma}=0$ does not cover the remaining cases. We give an example in Theorem 5.2.2 where $\mathrm{logTev}^{X_{r,s,a}}_{\Gamma}$ can be proven neither to be 0 nor the value conjectured in [CIL24, Conjecture 15].
@@ 5.1. Naive log quasimaps
Let $X=\mathrm{Bl}_{[0:1:0],[0:0:1]}({\mathbb{P}}^{2})$. A point of $X$ is given in coordinates by
@@
The pairs of coordinates $\{y_{1},y_{2}\},\{y_{2},y_{4}\},\{y_{3},y_{4}\},\{y_{1},y_{5}\},\{y_{3},y_{5}\}$ are not allowed to vanish simultaneously. The coordinates are taken up to simultaneous scaling of the collections of coordinates $\{y_{3},y_{4},y_{5}\},\{y_{1},y_{5}\},\{y_{2},y_{4}\}$. The 2-dimensional torus acts by scaling on $y_{4}$ and $y_{5}$. Let $D_{j}=\text{div}(y_{j})$ be the toric divisors. We identify the set $\Sigma(1)$ with $\{1,2,3,4,5\}$.
Fix tangency data $\Gamma$ for log maps to $f:{\mathbb{P}}^{1}\to X$. Let $d=\deg(f^{*}\mathcal{O}_{{\mathbb{P}}^{2}}(1))$. We construct a moduli space of naive log quasimaps as follows. Let $\mathcal{B}$ be the product of copies ${\mathbb{P}}^{1}_{j,v}\cong{\mathbb{P}}^{1}$ parametrizing points of the divisors $D_{j}$. As in §2.3, let $\nu:{\mathbb{P}}^{1}\times\mathcal{B}\to\mathcal{B}$ be the projection, and define the universal divisors $\mathcal{D}_{j,v},\mathcal{D}_{j}\subset{\mathbb{P}}^{1}\times\mathcal{B}$. Define the moduli space of naive log quasimaps
@@
taken up to scaling, and not all zero. Our space of naive log quasimaps $Q_{\Gamma}(X)$ is in some sense even more naive than the space $Q_{\Gamma}(X_{r,s,a})$ constructed in §2.3. This is because $(g_{3},g_{4},g_{5})$ in general do not even give rise to a quasimap in the sense of [CFK10], as we allow $(g_{3},g_{4})=(0,0)$ or $(g_{3},g_{5})=(0,0)$ as long as not all of $g_{3},g_{4},g_{5}$ vanish. However, this moduli space will suffice for our purposes.
We have a stratification by open subsets
@@
as in the case of $X_{r,s,a}$. The open subset $Q^{\neq 0}_{\Gamma}(X)$ is the locus where none of $g_{3},g_{4},g_{5}$ is identically zero. The moduli space of log maps $\mathcal{M}_{\Gamma}(X)$ is identified with the open subset of $Q^{\neq 0}_{\Gamma}(X)$ where none of the underlying divisors $D_{j}$ (where $j=1,\ldots,5$) share common points in their supports. Indeed, $g_{3},g_{4},g_{5}$ give rise to a map
(9) $f=[1_{D_{1}}1_{D_{2}}g_{3}:1_{D_{1}}g_{4}:1_{D_{2}}g_{5}]$
where $1_{D_{j}}$ is a non-zero element of $H^{0}({\mathbb{P}}^{1},\mathcal{O}(D_{j}))$. The sections $1_{D_{j}}$ can be chosen consistently over $\mathcal{B}$, by restricting a choice of global section of $1_{\mathcal{D}_{j}}$. The intermediate open set $\mathcal{M}^{\text{bpf}}_{\Gamma}(X)$ is the locus where the pairs of divisors $\{D_{1},D_{2}\},\{D_{2},D_{4}\},\{D_{3},D_{4}\},\{D_{1},D_{5}\},\{D_{3},D_{5}\}$ share no common points in their supports. A point $g=(g_{3},g_{4},g_{5})\in\mathcal{M}^{\text{bpf}}_{\Gamma}(X)$ also gives rise to a map $f:{\mathbb{P}}^{1}\to X$ via (9).
@@
$[V(p,x)]=(\zeta-\mathrm{H}_{1})(\zeta-\mathrm{H}_{2}).$

### v3

@@ 2.4. Incidence loci
on ${\mathbb{P}}(\mathcal{E}_{1})$. The first map is the inclusion of the tautological sub-line bundle on ${\mathbb{P}}(\mathcal{E}_{1})$. At a point of ${\mathbb{P}}(\mathcal{E}_{1})$, its image is the tuple of sections
$(g_{1},\ldots,g_{r+1})\in\bigoplus_{j=1}^{r+1}H^{0}({\mathbb{P}}^{1},\mathcal{O}(b)(-D_{j})).$
The second map, given an index $j\in\{2,\ldots,r+1\}$, returns the restriction of $x_{1}\cdot g_{j}-x_{j}\cdot g_{1}$ to $p\in{\mathbb{P}}^{1}$. It is induced by the inclusion $\mathcal{O}_{{\mathbb{P}}^{1}}(b)(-\mathcal{D}_{j})\subset\mathcal{O}_{{\mathbb{P}}^{1}}(b)$ (which amounts to viewing the sections $g_{j}$ instead as sections of $\mathcal{O}_{{\mathbb{P}}^{1}}(b)$) and the restriction to $p$.
Similarly, the equations (2) are the degeneracy loci of maps of line bundles
@@
where $\mathcal{O}_{Q}(-1)$ denotes the universal sub-line bundle on $Q_{\Gamma}(X_{r,s,a})$, and the first map is again the tautological inclusion. At a point of $Q_{\Gamma}(X_{r,s,a})$, its image is the tuple of sections
$(g_{r+2},\ldots,g_{r+s+1},g_{0})\in\left(\bigoplus_{j=r+2}^{r+s+1}H^{0}({\mathbb{P}}^{1},\mathcal{O}(c)(-D_{j}))\right)\oplus H^{0}({\mathbb{P}}^{1},\mathcal{O}(c-ab)(-D_{0})).$
The second map, given an index $j\in\{r+2,\ldots,r+s+1\}$, returns the restriction of $x_{j}\cdot g_{1}^{a}g_{0}-x_{1}^{a}x_{0}\cdot g_{j}$ to $p\in{\mathbb{P}}^{1}$. In particular, the restriction of second map to the last summand factors through the map
$\nu_{*}(\mathcal{O}_{{\mathbb{P}}^{1}}(c-ab)(-\mathcal{D}_{0}))\otimes\mathcal{O}_{{\mathbb{P}}(\mathcal{E}_{1})}(-a)\to\nu_{*}(\mathcal{O}_{{\mathbb{P}}^{1}}(c-ab))\otimes(\nu_{*}(\mathcal{O}_{{\mathbb{P}}^{1}}(b)))^{\otimes a}\to\nu_{*}(\mathcal{O}_{{\mathbb{P}}^{1}}(c))$
where the map $\mathcal{O}_{{\mathbb{P}}(\mathcal{E}_{1})}(-a)\to(\nu_{*}(\mathcal{O}_{{\mathbb{P}}^{1}}(b)))^{\otimes a}$ is induced from the universal sub-line bundle on ${\mathbb{P}}(\mathcal{E}_{1})$, and has image $g_{1}^{a}$ at a given point of $Q_{\Gamma}(X_{r,s,a})$.
@@ 3. Transversality
is of expected dimension 0. The degree of $V$ (as a 0-cycle) is therefore a candidate for the value of $\mathrm{logTev}^{X_{r,s,a}}_{\Gamma}$. In order to guarantee that $\deg(V)=\mathrm{logTev}^{X_{r,s,a}}_{\Gamma}$ (that is, the intersection number $\deg(V)$ is enumerative), we want $V$ to be reduced of dimension 0, and in addition supported in the locus $\mathcal{M}_{\Gamma}(X)\subset Q_{\Gamma}(X)$. Indeed, if this is the case, then $V$ is identified scheme-theoretically with a general fiber of the morphism $\tau:\mathcal{M}_{\Gamma,n}(X)\to\mathcal{M}_{0,n}\times X^{n}$.
@@ 3.1. Transversality for maps
where the last map is projection. If $n=\frac{m}{r+s}+1$, then this composition is a map between smooth varieties of the same dimension, so the first part of (i) follows from generic smoothness [hartshorne, Corollary 10.7]. To see that
@@ 3.2. Twisting
Denote by $f^{\prime}$ the value of $f$ after these steps are complete, and by $\Gamma^{\prime}$ the new tangency data. Output $f^{\prime}\in\mathcal{M}^{\text{bpf}}_{\Gamma^{\prime}}(X_{r,s,a})$.
@@
The first inequality is an equality if and only if no twisting is applied at points $q\neq p_{i}$, and the second inequality is an equality if and only if no twisting is applied at the $p_{i}$.
However, because the given $p_{i},x_{i}$ are chosen generally, $\mathcal{V}^{\prime}$ must dominate $({\mathbb{P}}^{1})^{n}\times X^{n}$. Indeed, the assumption is that there is at least one $f^{\prime}\in\mathcal{M}_{\Gamma^{\prime}}(X_{r,s,a})_{\Delta}$ satisfying the needed incidence conditions at a general collection of points $p_{i},x_{i}$. In particular, we need $\dim(\mathcal{V}^{\prime})\geq\dim(({\mathbb{P}}^{1})^{n}\times X^{n})$. Therefore, equality must hold everywhere. This means that no twisting was applied at all, which is to say that $f\in\mathcal{M}^{\text{bpf}}_{\Gamma}(X_{r,s,a})$. ∎
@@ 3.3. Vanishing sections
Combining with the fact that $m_{0}\geq\ell$ and m=(n1)(r+s)m=(n-1)(r+s) gives
@@ 4.1. Vanishing
which is equivalent to $m_{\ell}\leq n-1$. If instead $r=1$, then for $\ell=1,2$, the $(s+1)$-tuple of sections $g_{0}g_{\ell}^{a}:g_{3}:\cdots:g_{s+2}]$ defines a (log) map to ${\mathbb{P}}^{1}\to{\mathbb{P}}^{s}$ satisfying $n$ incidence conditions at general points. We find similarly, repeating the proof of Proposition 3.1.1 for $X={\mathbb{P}}^{s}$, that $m_{\ell}\leq n-1$.
Next, consider $\ell\in\{r+2,\ldots,r+s+1\}$. If $s>1$, then the sections $g_{0},\ldots,\widehat{g_{\ell}}\ldots,g_{r+s+1}$ define a point of $f^{\prime}\in\mathcal{M}_{\Gamma^{\prime}}(X_{r,s-1,a})$ satisfying $n$ general incidence conditions. Arguing as above gives $m_{\ell}\leq n-1$ once more. If $s=1$, then we consider instead the map $g:{\mathbb{P}}^{1}\to{\mathbb{P}}^{r}$ defined by $g_{1}:\cdots:g_{r+1}]$. We obtain in this case the stronger inequality $m_{0}+m_{r+2}\leq n-1$.
Finally, we prove the inequality $m_{r+2}+\cdots+m_{r+s+1}\geq(s-1)(n-1)$. There is nothing to check if $s=1$, so we assume that $s>1$, and consider the map $h:{\mathbb{P}}^{1}\to{\mathbb{P}}^{s-1}$ defined by $g_{r+2}:\cdots:g_{r+s+1}]$, which satisfies $n$ general incidence conditions. Therefore, we must have
@@ 4.2. Integral computation
parametrizes choices of the divisors $D_{j}$, and $\nu:{\mathbb{P}}^{1}\times\mathcal{B}\longrightarrow\mathcal{B}$ is the projection. The universal divisors are denoted $\mathcal{D}_{j}\subset{\mathbb{P}}^{1}\times\mathcal{B}$. We have the rank $r+1$ vector bundle
@@
For any $p\in{\mathbb{P}}^{1}$ and $x\in X^{\circ}$, recall from §2.4 that $V(p,x)$ is cut out by the vanishing of $r$ maps of line bundles
$\mathcal{O}_{{\mathbb{P}}(\mathcal{E}_{1})}(-1)\to\nu_{*}(\mathcal{O}_{{\mathbb{P}}^{1}}(b)|_{p})$
and $s$ maps of line bundles
$\mathcal{O}_{Q}(-1)\to\nu_{*}(\mathcal{O}_{{\mathbb{P}}^{1}}(c)|_{p}).$
Therefore, we have
@@
$=\zeta_{1}^{r}\zeta_{2}^{s},$
because the line bundles $\nu_{*}(\mathcal{O}_{{\mathbb{P}}^{1}}(b)|_{p})$ and $\nu_{*}(\mathcal{O}_{{\mathbb{P}}^{1}}(c)|_{p})$ are trivial, and
$-c_{1}(\mathcal{O}_{{\mathbb{P}}(\mathcal{E}_{1})}(-1))$ $=c_{1}(\mathcal{O}_{{\mathbb{P}}(\mathcal{E}_{1})}(1))=\zeta_{1},$
$-c_{1}(\mathcal{O}_{{\mathbb{P}}(\mathcal{E}_{2})}(-1))$ $=c_{1}(\mathcal{O}_{{\mathbb{P}}(\mathcal{E}_{1})}(1))=\zeta_{2}$
by definition.
@@
$\int_{{\mathbb{P}}(\mathcal{E}_{2})}\prod_{i=1}^{n}[V(p_{i},x_{i})]^{n}$ $=\int_{\mathcal{B}}\sum_{k_{0}+k_{r+2}+\cdots k_{r+s+1}=s(n-1)}\sum_{\ell=0}^{k_{0}}\binom{k_{0}}{\ell}a^{k_{0}-\ell}\mathcal{S}_{r(n-1)+k_{0}-\ell}(\mathcal{E}_{1})\mathrm{H}_{0}^{\ell}\mathrm{H}_{r+2}^{k_{r+2}}\cdots\mathrm{H}_{r+s+1}^{k_{r+s+1}}$
$=\int_{\mathcal{B}}\sum_{k_{0}+k_{r+2}+\cdots k_{r+s+1}=s(n-1)}\sum_{\ell=0}^{k_{0}}\binom{k_{0}}{\ell}a^{k_{0}-\ell}\sum_{k_{1}+\cdots k_{r+1}=r(n-1)+k_{0}-\ell}\mathrm{H}_{0}^{\ell}\prod_{j=1}^{r+s+1}\mathrm{H}_{j}^{k_{j}}.$
@@
because $n=\frac{m}{r+s}+1$. Note that this is only valid if $\ell=m_{0}\leq k_{0}$, because $\ell$ ranges between 0 and $k_{0}$ in the integrand. Thus, if $m_{0}>k_{0}$, then the integral is zero.
@@ 4.3. Specializations
Theorem 4.2.1 recovers [cil, Theorem 7] in the case of Hirzebruch surfaces, $r=s=1$. Our $m_{0},m_{1},m_{2},m_{3}$ are the $|\mu_{2}|,|\mu_{1}|,|\mu_{3}|,|\mu_{4}|$, respectively, appearing in [cil, Theorem 7]. If $s=1$ and $r\geq 1$ is arbitrary, then Theorem 4.2.1 proves [cil, Conjecture 14], and also makes precise exactly when the logarithmic Tevelev degree is zero. Taking $s$ arbitrary gives a more general statement.
The variety $X_{r,s,a}$ is Fano if and only if $a\leq s$. When in addition $a\geq 2$, the fixed-domain Gromov-Witten invariants of $X_{r,s,a}$ are shown not to be enumerative in [bllrst, §3.2], even in genus 0 for curve classes of large anti-canonical degree, disproving the main conjecture of [lp]. In contrast, the corresponding logarithmic invariants $\mathrm{logTev}^{X_{r,s,a}}_{\Gamma}$, with $\mu_{j}=(1)^{m_{j}}$, are enumerative in genus 0, and are determined by Theorem 4.2.1. The logarithmic setting is broadly understood to be more likely to yield enumerative counts of curves.
When $a=1$, we have $X_{r,s,a}\cong\mathrm{Bl}_{{\mathbb{P}}^{s-1}}{\mathbb{P}}^{r+s}$. Taking all of the $m_{j,v}$ equal to 1, substituting $c=d$ and b=dkb=d-k, and dividing by the factors of $m_{j}!$ that count orderings of the intersection points with the boundary gives that
@@
When $s=1$, this recovers the $\ell=1$ case of [cl2, Theorem 12] for the class $\beta$ of degree $d$ curves (against the hyperplane class) meeting the exceptional divisor with degree $k$. Without the logarithmic conditions, the $s>1$ case is new, but should also be no more difficult (even in higher genus, with some additional positivity assumptions on $\beta$) using the same methods of [cl2].
@@
Without tangency conditions, this specializes simply to $\mathrm{Tev}^{{\mathbb{P}}^{r}}_{0,n,d}=1$ [fl, Theorem 1.1].
@@ 5. Blow-ups of ${\mathbb{P}}^{2}$
The value of $\mathrm{logTev}^{X}_{\Gamma}$ when $X$ is the blow-up of ${\mathbb{P}}^{r}$ at $r$ points has also been predicted in [cil, Conjecture 15]. In this section, we show that the prediction does not always hold. For notational simplicity, we take $r=2$, though the arguments hold more generally.
@@
This is similar to the situation of $\mathrm{logTev}^{X_{r,s,a}}_{\Gamma}$, where we were able to show that, in the remaining cases, the logarithmic Tevelev degree vanishes. In the setting of blow-ups of ${\mathbb{P}}^{r}$, the range in which $\mathrm{logTev}^{X}_{\Gamma}=0$ does not cover the remaining cases. We give an example in Theorem 5.2.2 where $\mathrm{logTev}^{X_{r,s,a}}_{\Gamma}$ can be proven neither to be 0 nor the value conjectured in [cil, Conjecture 15].
@@ 5.1. Naive log quasimaps
Let $X=\mathrm{Bl}_{[0:1:0],[0:0:1]}({\mathbb{P}}^{2})$. A fan $\Sigma$ for $X$ is given below.
$\rho_{1}$$\rho_{2}$$\rho_{3}$$\rho_{4}$$\rho_{5}$
Figure 1. Fan of $X=\mathrm{Bl}_{[0:1:0],[0:0:1]}({\mathbb{P}}^{2})$
A point of $X$ is given in coordinates by
@@
The torus-invariant divisors $D_{j}=\text{div}(y_{j})$ given by the vanishing of $y_{j}$ correspond to the rays $\rho_{j}$ of the fan $\Sigma$ in Figure 1. We identify the set $\Sigma(1)$ with $\{1,2,3,4,5\}$. The pairs of coordinates $\{y_{1},y_{2}\},\{y_{2},y_{4}\},\{y_{3},y_{4}\},\{y_{1},y_{5}\},\{y_{3},y_{5}\}$ are not allowed to vanish simultaneously. The coordinates are taken up to simultaneous scaling of the collections of coordinates $\{y_{3},y_{4},y_{5}\},\{y_{1},y_{5}\},\{y_{2},y_{4}\}$. The 2-dimensional torus acts by scaling on $y_{4}$ and $y_{5}$.
Fix tangency data $\Gamma$ for log maps $f:{\mathbb{P}}^{1}\to X$. Let $d=\deg(f^{*}\mathcal{O}_{{\mathbb{P}}^{2}}(1))$. We construct a moduli space of naive log quasimaps as follows. Let $\mathcal{B}$ be the product of copies ${\mathbb{P}}^{1}_{j,v}\cong{\mathbb{P}}^{1}$ parametrizing points of the divisors $D_{j}$. As in §2.3, let $\nu:{\mathbb{P}}^{1}\times\mathcal{B}\to\mathcal{B}$ be the projection, and define the universal divisors $\mathcal{D}_{j,v},\mathcal{D}_{j}\subset{\mathbb{P}}^{1}\times\mathcal{B}$. Define the moduli space of naive log quasimaps
@@
taken up to scaling, and not all zero.
A general point of $Q_{\Gamma}(X)$ gives rise to a log map $f:{\mathbb{P}}^{1}\to X$. Indeed, suppose that none of $g_{3},g_{4},g_{5}$ is identically zero and that none of the underlying divisors $D_{j}$ (where $j=1,\ldots,5$) share common points in their supports. For $j=1,2$, let $1_{D_{j}}$ be a non-zero element of $H^{0}({\mathbb{P}}^{1},\mathcal{O}(D_{j}))$. Then,
(9) $f=[1_{D_{1}}1_{D_{2}}g_{3}:1_{D_{1}}g_{4}:1_{D_{2}}g_{5}]$
is a point of $\mathcal{M}_{\Gamma}(X)$. Moreover, the sections $1_{D_{j}}$ can be chosen consistently over $\mathcal{B}$, by restricting a choice of global section of $1_{\mathcal{D}_{j}}$. In this way, $\mathcal{M}_{\Gamma}(X)$ is identified with an open subset of $Q_{\Gamma}(X)$.
Let $Q^{\neq 0}_{\Gamma}(X)\subset Q_{\Gamma}(X)$ be the open subset where none of $g_{3},g_{4},g_{5}$ is identically zero. Let $\mathcal{M}^{\text{bpf}}_{\Gamma}(X)\subset Q^{\neq 0}_{\Gamma}(X)$ be the intermediate open set where the pairs of divisors
$\{D_{1},D_{2}\},\{D_{2},D_{4}\},\{D_{3},D_{4}\},\{D_{1},D_{5}\},\{D_{3},D_{5}\}$
share no common points in their supports. A point $g=(g_{3},g_{4},g_{5})\in\mathcal{M}^{\text{bpf}}_{\Gamma}(X)$ also gives rise to a map $f:{\mathbb{P}}^{1}\to X$ via (9). We have a stratification by open subsets
@@
as in the case of $X_{r,s,a}$.
Our space of naive log quasimaps $Q_{\Gamma}(X)$ is in some sense even more naive than the space $Q_{\Gamma}(X_{r,s,a})$ constructed in §2.3. This is because $(g_{3},g_{4},g_{5})$ in general do not even give rise to a quasimap in the sense of [cfk], as we allow $(g_{3},g_{4})=(0,0)$ or $(g_{3},g_{5})=(0,0)$ as long as not all of $g_{3},g_{4},g_{5}$ vanish. However, this moduli space will suffice for our purposes.
@@
$V(p,x)]=(\zeta-\mathrm{H}_{1})(\zeta-\mathrm{H}_{2}).$